hooks law states that the distance d that a spring is stretched by a hangin object varies directly as the weight w of the object. if a spring is stretched to 40 cm by a 3-kg barbell, what is the distance stretched by a 5-kg barbell?
Answer by stanbon(75887)

hooks law states that the distance d that a spring is stretched by a hanging object varies directly as the weight w of the object.
------------------------
if a spring is stretched to 40 cm by a 3-kg barbell, what is the distance stretched by a 5-kg barbell?
--------------------------
d = kw
---
Solve for "k" using "a spring is stretched to 40 cm by a 3-kg barbell":
40 = k*3
k = 40/3
---------------
Equation:
d = (40/3)w
---
Solve:what is the distance stretched by a 5-kg barbell?
d = (40/3)*5
---
d = 200/3
distance = 66 2/3 cm
==================
Cheers,
Stan H.